OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 3x - 4y = -12.

OpenStudy (anonymous):

x/(-12/3)+y/(-12/-4)=1 x/(-4)+y/(3)=1 x int = -4, y int =3

hero (hero):

3(0) - 4y = -12 y = 3 3x - 4(0) = -12 x = -4